### v3

# Akchen and Mišić

[…]
Akchen and Mišić
[…]
Assortment Optimization under the Decision Forest Model
[…]
Assortment Optimization under the Decision Forest Model
[…]
decision trees; choice modeling; integer optimization; Benders decomposition; choice overload
1 Introduction
Assortment optimization is a basic operational problem faced by many firms. In its simplest form, the problem can be posed as follows. A firm has a set of products that it can offer, and a set of customers who have preferences over those products; what is the set of products the firm should offer so as to maximize the revenue that results when the customers choose from these products?
While assortment optimization and the related problem of product line design have been studied extensively under a wide range of choice models, the majority of research in this area focuses on rational choice models, specifically those that follow the random utility maximization (RUM) assumption. A significant body of research in the behavioral sciences shows that customers behave in ways that deviate significantly from predictions made by RUM models. In addition, there is a substantial number of empirical examples of firms that make assortment decisions in ways that directly exploit customer irrationality. For example, the paper of Kivetz et al. (2004) provides an example of an assortment of document preparation systems from Xerox that are structured around the decoy effect, and an example of an assortment of servers from IBM that are structured around the compromise effect.
A recent paper by Chen and Mišić (2022) proposed a new choice model called the decision forest model for capturing customer irrationalities. This model involves representing the customer population as a probability distribution over binary trees, with each tree representing the decision process of one customer type. In a key result of the paper, the authors showed that this model is universal: every discrete choice model is representable as a decision forest model. While the paper of Chen and Mišić (2022) alludes to the downstream assortment optimization problem, it is entirely focused on model representation and prediction: it does not provide any answer to how one can select an optimal assortment with respect to a decision forest model.
In the present paper, we present a methodology for assortment optimization under the decision forest model, based on mixed-integer optimization. Our approach allows the firm to obtain assortments that are either provably optimal or within a desired optimality gap for a given decision forest model. At the same time, the approach easily allows the firm to incorporate business rules as linear constraints in the MIO formulation. Most importantly, given the universality property of the decision forest model, our optimization approach allows a firm to optimally tailor its assortment to any kind of predictable irrationality in the firm’s customer population.
We make the following specific contributions:
  1. 1.
    We formally define the assortment optimization problem under the decision forest model and analyze its computational hardness. Specifically, assuming the tree depth is bounded by a constant $k$, we prove that it is NP-hard to approximate the problem within a factor that grows exponentially with $k$. Given this computational challenge, we then adopt an integer programming approach and present two formulations: SplitMIO and ProductMIO. We analyze the constraint matrices of the two formulations and show that in the special case of a single purchase decision tree, SplitMIO is integral when each product appears at most once in the splits of the tree, and ProductMIO is always integral regardless of the tree structure.
  2. 2.
    We propose a Benders decomposition approach for solving the two formulations at a large scale. We show that Benders cuts for the linear optimization relaxations of SplitMIO can be obtained via a greedy algorithm that solves both the primal and dual of the subproblem. We also provide a simple example to show that the same type of greedy algorithm fails to solve the primal subproblem of ProductMIO. We further show how to obtain Benders cuts for the integer solutions of both SplitMIO and ProductMIO in closed form.
  3. 3.
    We present numerical experiments using both synthetic and real-world data. We first use the synthetic data to examine the formulation strength of SplitMIO and ProductMIO models, and to demonstrate the scalability of the Benders decomposition approach for the SplitMIO formulation in problem instances involving up to 3000 products, 500 trees, and 512 leaves per tree. We then use a real-world dataset to demonstrate how the decision forest model can factor a behavioral anomaly (choice overload) into its assortment decision and create higher revenue.
We organize the paper as follows. Section 2 reviews the related literature. Section 3 defines the assortment problem, characterizes its inapproximability, and presents the two MIO formulations. Section 4 proposes a Benders decomposition approach for these formulations. Sections 5 and 6 present the numerical results involving both the synthetic and real-world data. All proofs are relegated to the appendix.
2 Literature review
Assortment optimization has been extensively studied in the operations management community; we refer readers to Gallego and Topaloglu (2019) for a recent review of the literature. The literature on assortment optimization has focused on developing approaches for finding the optimal assortment under many different rational choice models, such as the multinomial logit model (MNL) model (Talluri and Van Ryzin 2004, Sumida et al. 2020), the latent class MNL model (Bront et al. 2009, Méndez-Díaz et al. 2014), the Markov chain choice model (Feldman and Topaloglu 2017, Désir et al. 2020) and the ranking-based model (Aouad et al. 2020, 2018, Feldman et al. 2019).
In addition to the assortment optimization literature, our paper is also related to the literature on product line design found in the marketing community. While assortment optimization is more often focused on the tactical decision of selecting which existing products to offer, where the products are ones that have been sold in the past and the choice model comes from transaction data involving those products, the product line design problem involves selecting which new products to offer, where the products are candidate products (i.e., they have not been offered before) and the choice model comes from conjoint survey data, where customers are asked to rate or choose between hypothetical products. Research in this area has considered different approaches to solve the problem under the ranking-based/first-choice model (McBride and Zufryden 1988, Belloni et al. 2008, Bertsimas and Mišić 2019) and the MNL model (Chen and Hausman 2000, Schön 2010).
Our paper is related to Bertsimas and Mišić (2019), which presents integer optimization formulations of the product line design problem when the choice model is a ranking-based model. As we will see later, our formulation SplitMIO can be viewed a generalization of the formulation Bertsimas and Mišić (2019), to the decision forest model. In addition, the paper of Bertsimas and Mišić (2019) develops a specialized Benders decomposition approach for its formulation, which uses the fact that one can solve the subproblem associated with each customer type by applying a greedy algorithm. We will show in Section 4 that this same property generalizes to the SplitMIO formulation, leading to a tailored Benders decomposition algorithm for solving SplitMIO at scale.
Beyond these specific connections, the majority of the literature on assortment optimization and product line design considers rational choice models, whereas our paper contributes a methodology for non-rational assortment optimization. Fewer papers have focused on choice modeling for non-rational customer behavior; besides the decision forest model, other models include the generalized attraction model (GAM; Gallego et al. 2015), the generalized stochastic preference model (Berbeglia 2018) and the generalized Luce model (Echenique and Saito 2019). An even smaller set of papers has considered assortment optimization under non-rational choice models, which we now review. The paper of Flores et al. (2017) considers assortment optimization under the two-stage Luce model, and develops a polynomial time algorithm for solving the unconstrained assortment optimization problem. The paper of Rooderkerk et al. (2011) considers a context-dependent utility model where the utility of a product can depend on other products that are offered and that can capture compromise, attraction and similarity effects; the paper empirically demonstrates how incorporating context effects leads to a predicted increase of 5.4% in expected profit.
Relative to these papers, our paper differs in that it considers the decision forest model. As noted earlier, the decision forest model can represent any type of choice behavior, and as such, an assortment optimization methodology based on such a model is attractive in terms of allowing a firm to take the next step from a high-fidelity model to a decision. In addition, our methodology is built on mixed-integer optimization. This is advantageous because it allows a firm to leverage continuing improvements in integer optimization solvers such as Gurobi (Gurobi Optimization, LLC 2024) and CPLEX (IBM 2024), as well as continuing improvements in computer hardware. At the same time, integer optimization allows firms to accommodate business requirements using linear constraints, enhancing the practical applicability of the approach. Lastly, integer optimization also allows one to take advantage of well-studied large-scale solution methods. One such method that we focus on in this paper is Benders decomposition, which has seen an impressive resurgence in recent years for delivering state-of-the-art performance on large-scale problems such as hub location (Contreras et al. 2011), facility location (Fischetti et al. 2017) and set covering (Cordeau et al. 2019); see also Rahmaniani et al. (2017) for a review of the recent literature.
In addition to the assortment optimization and product line design literatures, our formulations also have connections with others that have been proposed in the broader optimization literature. The formulation SplitMIO that we will present later can be viewed as a special case of the mixed-integer optimization formulation of Mišić (2020) for optimizing the predicted value of a tree ensemble model, such as a random forest or a boosted tree model. The other formulation, ProductMIO, also has a connection to the integer optimization literature on formulating disjunctive constraints through independent branching schemes (Vielma et al. 2010, Vielma and Nemhauser 2011, Huchette and Vielma 2019); we also discuss this connection in more detail in Section 3.4.
3 Optimization model
In this section, we define the decision forest assortment optimization problem (Section 3.1), analyze its computational complexity (Section 3.2), and subsequently present the two mixed-integer formulations, SplitMIO (Section 3.3) and ProductMIO (Section 3.4). Lastly, we show that when the number of trees is assumed to be constant, the decision forest assortment optimization problem can be solved in polynomial time (Section 3.5).
3.1 Problem definition
We first briefly review the decision forest model of Chen and Mišić (2022), and then formally state the assortment optimization problem. We assume that there are $n$ products, indexed from 1 to $n$, and let $\mathcal{N}=\{1,\dots,n\}$ denote the set of all products. An assortment $S$ corresponds to a subset of $\mathcal{N}$. When offered $S$, a customer may choose to purchase one of the products in $S$, or to not purchase anything from $S$ at all; we use the index 00 to denote the latter possibility, which we will also refer to as the no-purchase option or the outside option.
The basic building block of the decision forest model is a purchase decision tree. Each tree represents the purchasing behavior of one type of customer when presented with an assortment $S$. Specifically, for an assortment $S$, the customer behaves as follows: the customer starts at the root of the tree. The customer checks whether the product corresponding to the root node is contained in $S$; if it is, she proceeds to the left child, and if not, she proceeds to the right child. She then checks again with the product at the new node, and the process repeats, until the customer reaches a leaf; the option that is at the leaf represents the choice of that customer. Figure 1 (left) shows an example of a purchase decision tree being used to map an assortment to a purchase decision.
Figure 1: (left) Example of a purchase decision tree for $n=5$ products. Leaf nodes are enclosed in squares, while split nodes are not enclosed. The number on each node corresponds either to $v(t,s)$ for splits, or $c(t,\ell)$ for leaves. The path highlighted in red indicates how a customer following this tree to maps the assortment $S=\{1,3,4,5\}$ to a leaf. For this assortment, the customer’s decision is to purchase product 5; (right) Example of a purchase decision tree that represents a preference ranking $\{3\succ 1\succ 2\succ 0\}$.
Formally, a purchase decision tree is a directed binary tree, with each leaf node corresponding to an option in $\mathcal{N}\cup\{0\}$, and each non-leaf (or split) node corresponding to a product in $\mathcal{N}$. We use $\mathbf{splits}(t)$ to denote the set of split nodes of tree $t$, and $\mathbf{leaves}(t)$ to denote the set of leaf nodes. We use $c(t,\ell)$ to denote the purchase decision of leaf $\ell$ of tree $t$, i.e., the option chosen by tree $t$ if the assortment is mapped to leaf $\ell$. We use $v(t,s)$ to denote the product that is checked at split node $s$ in tree $t$. We make the following assumption about our purchase decision trees. {assumption} Let $t$ be a purchase decision tree. For any two split nodes $s$ and $s^{\prime}$ of $t$ such that $s^{\prime}$ is a descendant of $s$, $v(t,s)\neq v(t,s^{\prime})$. This assumption states that once a product appears on a split $s$, it cannot appear on any subsequent split $s^{\prime}$ that is reached by proceeding to the left or right child of $s$; in other words, each product in $\mathcal{N}$ appears at most once along the path from the root node to a leaf node, for every leaf node. As discussed in Chen and Mišić (2022), this assumption is not restrictive, as any tree for which this assumption is violated has a set of splits and leaves that are redundant and unreachable, and the tree can be modified to obtain an equivalent tree that satisfies the assumption.
The decision forest model assumes that the customer population is represented by a collection of trees or a forest $F$. Each tree $t\in F$ corresponds to a different customer type. We use $\lambda_{t}$ to denote the probability associated with customer type/tree $t$, and $\boldsymbol{\lambda}=(\lambda_{t})_{t\in F}$ to denote the probability distribution over the forest $F$. For each tree $t$, we use $\eta(t,S)$ to denote the choice that a customer type following tree $t$ will make when given the assortment $S$. For a given assortment $S\subseteq\mathcal{N}$ and a given choice $j\in S\cup\{0\}$, we use $\mathbf{P}^{(F,\boldsymbol{\lambda})}(j\mid S)$ to denote the choice probability, i.e., the probability of a random customer choosing $j$ when offered the assortment $S$. It is defined as
$\mathbf{P}^{(F,\boldsymbol{\lambda})}(j\mid S)=\sum_{t\in F}\lambda_{t}\cdot\mathbb{I}\{{\color[rgb]{0,0,0}\eta(t,S)}=j\}.$ (1)
We now define the assortment optimization problem. We use $\rho_{i}$ to denote the marginal revenue of product $i$; for convenience, we use $\rho_{0}=0$ to denote the revenue of the no-purchase option. The assortment optimization problem that we wish to solve is
$\underset{S\subseteq\mathcal{N}}{\text{maximize}}\ \sum_{i\in S}{\color[rgb]{0,0,0}\rho_{i}}\cdot\mathbf{P}^{(F,\boldsymbol{\lambda})}(i\mid S).$ (2)
This is a challenging problem due to the generality of the choice model $\mathbf{P}^{(F,\boldsymbol{\lambda})}(\cdot\mid\cdot)$. Notably, it subsumes the assortment optimization problem under the ranking-based model, which is known to be NP-hard (Aouad et al. 2018, Feldman et al. 2019). This conclusion arises from the observation that, in the special case where each tree $t\in F$ branches exclusively to the right – meaning the left child of every split node in tree $t$ is always a leaf node (as illustrated in the right panel of Figure 1) – the forest $F$ simplifies into a collection of preference rankings. This observation directly leads to the following corollary.
Corollary 3.1
The assortment optimization problem (2) is NP-Hard.
3.2 Inapproximability Results
We strengthen the complexity results for the assortment optimization problem (2) by providing an inapproximability gap based on the depth of the decision trees. First, we define the depth of a tree $t$ as the maximal distance from the root of tree $t$ to any leaf node $\ell\in\mathbf{leaves}(t)$. For example, both trees in Figure 1 are of depth three. Second, we let $\text{AODF}(k)$ denote problem (2) where each tree $t\in F$ can have depth at most $k$. With these definitions, we have the following result.
Theorem 3.2
If P $\neq$ NP, for every constant $k\geq 7$ and $\epsilon>0$, there is no polynomial-time $\left(2^{k-2\lceil\sqrt{k}\rceil}/(2ek)-\epsilon\right)$-approximate algorithm for AODF(k+1), even when each tree has equal probability weight in the model, i.e., $\lambda_{t}=1/|F|$.
[…]
Theorem 3.2 implies that the inapproximability gap of the assortment problem (2) grows at least exponentially with respect to the tree depth, since the ratio in Theorem 3.2 follows $2^{k-2\lceil\sqrt{k}\rceil}/(2ek)=2^{k-2\lceil\sqrt{k}\rceil-\log k-1-\log_{2}e}=2^{k-o(k)}$. The following proposition further shows that this exponential dependence is asymptotically tight in $k$ (i.e., when $k$ is large), as there exists an approximation algorithm achieving approximation ratio $2^{k}$.
Proposition 3.3
AODF(k) can be approximated within factor $2^{k}$.
Before we proceed to solve the challenging assortment optimization problem (2), we pause to make two remarks on Theorem 3.2. First, recall that any ranking of length $k$ can be represented by a decision tree that is constrained to branch to the right and of depth $k$. Aouad et al. (2018) shows that the assortment optimization under the ranking-based model is NP-hard to be approximated within $O(k^{1-\epsilon})$ factor provided that each ranking in the model has a length of at most $k$ (see Section 3.2 of Aouad et al. (2018)). Theorem 3.2 shows that the assortment problem becomes much harder when optimizing over decision trees instead of rankings. Specifically, the inapproximability factor grows from a linear rate $O(k)$ to an exponential rate $O(2^{k})$ as a function of $k$.
Second, together with Chen and Mišić (2022), Theorem 3.2 highlights a fundamental tradeoff between the representation power and tractability in the decision forest model. In particular, Chen and Mišić (2022) shows that as the decision trees grow deeper, the model becomes more expressive in capturing choice behavior; at its maximum depth, $n$, the decision forest model can represent any discrete choice model (see Theorems 1 and 2 in Chen and Mišić (2022)). At the same time, our Theorem 3.2 establishes that deeper trees also lead to increased computational complexity for assortment optimization, with inapproximability growing exponentially in tree depth. This tradeoff underscores the importance of managing tree depth to ensure that the model accurately reflects consumer choice while maintaining tractability for downstream operational tasks.
Given the inapproximability of problem (2), we consider a mixed-integer optimization (MIO) approach to solve the problem and present two formulations in the next two subsections.
3.3 Formulation 1: SplitMIO
We now present our first formulation of the assortment optimization problem (2) as a mixed-integer optimization (MIO) problem. To formulate the problem, we introduce some additional notation. For notational convenience we let $r_{t,\ell}=\rho_{c(t,\ell)}$ be the revenue of the purchase option of leaf $\ell$ of tree $t$. We let $\mathbf{left}(s)$ denote the set of leaf nodes that are to the left of split $s$ (i.e., can only be reached by taking the left branch at split $s$), and similarly, we let $\mathbf{right}(s)$ denote the leaf nodes that are to the right of $s$. We introduce two sets of decision variables. For each $i\in\mathcal{N}$, we let $x_{i}$ be a binary decision variable that is 1 if product $i$ is included in the assortment, and 0 otherwise. For each tree $t\in F$ and leaf $\ell\in\mathbf{leaves}(t)$, we let $y_{t,\ell}$ be a binary decision variable that is 1 if the assortment encoded by $\mathbf{x}$ is mapped to leaf $\ell$ of tree $t$, and 0 otherwise.
With these definitions, our first formulation, SplitMIO, is given below.
[…]
subject to $\sum_{\ell\in\mathbf{leaves}(t)}y_{t,\ell}=1,\quad\forall\ t\in F,$ (3b)
$\sum_{\ell\in\mathbf{left}(s)}y_{t,\ell}\leq x_{v(t,s)},\quad\forall\ t\in F,\ s\in\mathbf{splits}(t),$ (3c)
$\sum_{\ell\in\mathbf{right}(s)}y_{t,\ell}\leq 1-x_{v(t,s)},\quad\forall\ t\in F,\ s\in\mathbf{splits}(t),$ (3d)
$x_{i}\in\{0,1\},\quad\forall\ i\in\mathcal{N},$ (3e)
$y_{t,\ell}\geq 0,\quad\forall\ t\in F,\ \ell\in\mathbf{leaves}(t).$ (3f)
In order of appearance, the constraints in this formulation have the following meaning. Constraint (3b) requires that for each customer type $t$, the assortment encoded by $\mathbf{x}$ is mapped to exactly one leaf. Constraint (3c) imposes that for a split $s$ in tree $t$, if product $v(t,s)$ is not in the assortment, then the assortment cannot be mapped to any of the leaves that are to the left of split $s$ in tree $t$. Constraint (3d) is the symmetric case of constraint (3c) for the right-hand subtree of split $s$ of tree $t$. The last two constraints require that $\mathbf{x}$ is binary and $\mathbf{y}$ is nonnegative. Note that it is not necessary to require $\mathbf{y}$ to be binary, as the constraints ensure that each $y_{t,\ell}$ automatically takes the correct value whenever $\mathbf{x}$ is binary. Finally, the objective function corresponds to the expected per-customer revenue of the assortment.
The SplitMIO formulation is related to two existing MIO formulations in the literature. First, it can be viewed as a specialized case of the MIO formulation in Mišić (2020). In that paper, the author develops a formulation for tree ensemble optimization, i.e., the problem of setting the independent variables in a tree ensemble model (e.g., a random forest or a gradient boosted tree model) to maximize the value predicted by that ensemble. Later, in Section 4, we introduce a specialized algorithm that leverages a primal-dual greedy approach to efficiently solve the linear relaxation of SplitMIO. This method, which enhances the tractability of the SplitMIO formulation, represents our contribution to the literature and was not considered in Mišić (2020). Second, the SplitMIO formulation also relates to the MIO formulation for the product line design problem under the ranking-based model (Bertsimas and Mišić 2019). Recall that the ranking-based model can be regarded as a special case of the decision forest model (illustrated in the right panel of Figure 1). In this special case, the formulation (3) can be reduced to the MIO formulation for product line design under ranking-based models presented in Bertsimas and Mišić (2019).
Before continuing to our second formulation, we establish an important property of problem (3) when $|F|=1$. When $|F|=1$, we can show that $\mathcal{F}_{\textsc{SplitMIO}}$ is integral in a particular special case. (Note that in the statement of the proposition below, we drop the index $t$ to simplify notation.)
Proposition 3.4
Let $(F,\boldsymbol{\lambda})$ be a decision forest model consisting of a single tree, i.e., $|F|=1$. Then $\mathcal{F}_{\textsc{SplitMIO}}$ is integral, i.e., every extreme point $(\mathbf{x},\mathbf{y})$ of the polyhedron $\mathcal{F}_{\textsc{SplitMIO}}$ satisfies $\mathbf{x}\in\{0,1\}^{N}$, if and only if $v(s)=i$ for at most one $s\in\mathbf{splits}$ for every product $i\in\mathcal{N}$.
The proof of this result (see Appendix 8.3) follows by showing that the constraint matrix defining $\mathcal{F}_{\textsc{SplitMIO}}$ is totally unimodular. Proposition 3.4 is significant because it implies that for $|F|=1$, the distinction between trees where each product appears at most once in any split and trees where a product may appear two or more times as a split is sharp. This insight provides the motivation for our second formulation, ProductMIO, which we present next.
3.4 Formulation 2: ProductMIO
The second formulation of problem (2) that we will present is motivated by the behavior of SplitMIO when a product participates in two or more splits. In particular, observe that a product $i$ may participate in two different splits $s_{1}$ and $s_{2}$ in the same decision tree $t$. In this case, constraint (3c) in the SplitMIO will result in two constraints:
$\sum_{\ell\in\mathbf{left}(s_{1})}y_{t,\ell}\leq x_{i},$ (4)
$\sum_{\ell\in\mathbf{left}(s_{2})}y_{t,\ell}\leq x_{i}.$ (5)
In the above two constraints, observe that $\mathbf{left}(s_{1})$ and $\mathbf{left}(s_{2})$ are disjoint (this is a direct consequence of Assumption 1). Given this and constraint (3b) that requires the $y_{t,\ell}$ variables to sum to 1, we can come up with a constraint that strengthens constraints (4) and (5) by combining them:
$\sum_{\ell\in\mathbf{left}(s_{1})}y_{t,\ell}+\sum_{\ell\in\mathbf{left}(s_{2})}y_{t,\ell}\leq x_{i}.$ (6)
In general, one can aggregate all the $y_{t,\ell}$ variables that are to the left of all splits involving a product $i$ to produce a single left split constraint for product $i$. The same can also be done for the right split constraints. Generalizing this principle leads to the following alternate formulation, which we refer to as ProductMIO:
[…]
subject to $\sum_{\ell\in\mathbf{leaves}(t)}y_{t,\ell}=1,\quad\forall\ t\in F,$ (7b)
$\sum_{\begin{subarray}{c}s\in\mathbf{splits}(t):\\ v(t,s)=i\end{subarray}}\ \sum_{\ell\in\mathbf{left}(s)}y_{t,\ell}\leq x_{i},\quad\forall\ t\in F,\ i\in\mathcal{N},$ (7c)
$\sum_{\begin{subarray}{c}s\in\mathbf{splits}(t):\\ v(t,s)=i\end{subarray}}\ \sum_{\ell\in\mathbf{right}(s)}y_{t,\ell}\leq 1-x_{i},\quad\forall\ t\in F,\ i\in\mathcal{N},$ (7d)
$x_{i}\in\{0,1\},\quad\forall\ i\in\mathcal{N},$ (7e)
$y_{t,\ell}\geq 0,\quad\forall\ t\in F,\ \ell\in\mathbf{leaves}(t).$ (7f)
ProductMIO differs from SplitMIO in several ways. First, note that while both formulations have the same number of variables, formulation ProductMIO has a smaller number of constraints. In particular, SplitMIO has one left and one right split constraints for each split in each tree, whereas ProductMIO has one left and one right split constraint for each product. When the trees involve a large number of splits, this can lead to a sizable reduction in the number of constraints. Note also that when a product does not appear in any splits of a tree, we can also safely omit constraints (7c) and (7d) for that product.
The second difference with formulation SplitMIO, as we have already mentioned, is in formulation strength. Let $\mathcal{F}_{\textsc{ProductMIO}}$ be the feasible region of the linear optimization (LO) relaxation of ProductMIO. The following proposition formalizes the fact that formulation ProductMIO is at least as strong as formulation SplitMIO.
Proposition 3.5
For any decision forest model $(F,\boldsymbol{\lambda})$, $\mathcal{F}_{\textsc{ProductMIO}}\subseteq\mathcal{F}_{\textsc{SplitMIO}}$.
The proof follows straightforwardly using the logic given above; we thus omit the proof.
The last major difference is in how ProductMIO behaves when $|F|=1$. We saw that a sufficient condition for $\mathcal{F}_{\textsc{SplitMIO}}$ to be integral when $|F|=1$ is that each product appears in at most one split in the tree. In contrast, formulation ProductMIO is always integral when $|F|=1$.
Proposition 3.6
For any decision forest model $(F,\boldsymbol{\lambda})$ with $|F|=1$, $\mathcal{F}_{\textsc{ProductMIO}}$ is integral.
In contrast to Proposition 3.4, we take a different approach to proving the integrality property, as the constraint matrix of the ProductMIO formulation is generally not totally unimodular (see the example in Section 8.4). The proof of Proposition 3.6, given in Appendix 8.5, leverages a connection between ProductMIO and another type of formulation in the literature. In particular, a stream of papers in the mixed-integer optimization community (Vielma et al. 2010, Vielma and Nemhauser 2011, Huchette and Vielma 2019) has considered a general approach for deriving small and strong formulations of disjunctive constraints using independent branching schemes; we briefly review the most general such approach from Huchette and Vielma (2019) and showcase its connection to ProductMIO. In this approach, one has a finite ground set $J$, and is interested in optimizing over a particular subset of the $(|J|-1)-$dimensional unit simplex over $J$, $\Delta^{J}=\{\boldsymbol{\lambda}\in\mathbb{R}^{J}\mid\sum_{j\in J}\lambda_{j}=1;\boldsymbol{\lambda}\geq\mathbf{0}\}$. The specific subset that we are interested in is called a combinatorial disjunctive constraint (CDC), and is given by $\text{CDC}(\mathcal{S})=\bigcup_{S\in\mathcal{S}}Q(S)$, where $\mathcal{S}$ is a finite collection of subsets of $J$ and $Q(S)=\{\boldsymbol{\lambda}\in\Delta\mid\lambda_{j}\leq 0\ \text{for}\ j\in J\setminus S\}$ for any $S\subseteq J$. This approach is very general: for example, by associating each $j$ with a point $\mathbf{x}^{j}$ in $\mathbb{R}^{n}$, one can use $\text{CDC}(\mathcal{S})$ to model an optimization problem over a union of polyhedra, where each polyhedron is the convex hull of a collection of vertices in $S\in\mathcal{S}$.
A $k$-way independent branching scheme of depth $t$ is a representation of $\text{CDC}(\mathcal{S})$ as a sequence of $t$ choices between $k$ alternatives. Specifically, we have $\text{CDC}(\mathcal{S})=\bigcap_{m=1}^{t}\bigcup_{i=1}^{k}Q({\color[rgb]{0,0,0}L^{i}_{m}})$, where ${\color[rgb]{0,0,0}L^{i}_{m}}\subseteq J$. In the special case that $k=2$, we can write $\text{CDC}(\mathcal{S})=\cap_{m=1}^{t}(L_{{m}}\cup R_{{m}})$ where $L_{{m}},R_{{m}}\subseteq J$. This representation is known as a pairwise independent branching scheme and the constraints of the corresponding MIO can be written simply as
$\sum_{j\in L_{{m}}}\lambda_{j}\leq z_{m},\quad\forall\ m\in\{1,\dots,t\},$ (8a)
$\sum_{j\in R_{{m}}}\lambda_{j}\leq 1-z_{m},\quad\forall\ m\in\{1,\dots,t\},$ (8b)
$z_{m}\in\{0,1\},\quad\forall\ m\in\{1,\dots,t\},$ (8c)
$\sum_{j\in J}\lambda_{j}=1,$ (8d)
$\lambda_{j}\geq 0,\quad\forall\ j\in J.$ (8e)
This particular special case is important because it is always integral (see Theorem 1 of Vielma et al. 2010). Moreover, we can see that ProductMIO bears a strong resemblance to formulation (8). Constraints (8a) and (8a) correspond to constraints (7c) and (7d), respectively. In terms of variables, the $\lambda_{j}$ and $z_{m}$ variables in formulation (8) correspond to the $y_{t,\ell}$ and $x_{i}$ variables in ProductMIO, respectively.
One notable difference is that in practice, one would use formulation (8) in a modular way; specifically, one would be faced with a problem where the feasible region can be written as $\text{CDC}(\mathcal{S}_{1})\cap\text{CDC}(\mathcal{S}_{2})\cap\dots\cap\text{CDC}(\mathcal{S}_{G})$, where each $\mathcal{S}_{g}$ is a collection of subsets of $J$. To model this feasible region, one would introduce a set of $z_{g,m}$ variables for the $g$th CDC, enforce constraints (8a) - (8c) for the $g$th CDC, and use only one set of $\lambda_{j}$ variables for the whole formulation. Thus, the $\lambda_{j}$ variables are the “global” variables, while the $z_{g,m}$ variables would be “local” and specific to each CDC. In contrast, in ProductMIO, the $x_{i}$ variables (the analogues of $z_{m}$) are the “global” variables, while the $y_{t,\ell}$ variables (the analogues of $\lambda_{j}$) are the “local” variables.
3.5 Polynomial time algorithm when $|F|=O(1)$
We conclude Section 3 by discussing a special setting in which the assortment problem (2) is efficiently solvable. When $|F|=1$, the assortment optimization problem is trivial to solve, because one can simply find which leaf corresponds to the highest value of $r_{t,\ell}$, and then follow the path from the leaf to the root backwards to identify which products must be present and absent in the assortment to ensure that it is mapped to this leaf. If the number of leaves is polynomial in $n$, then one can solve this problem in time polynomial in $n$.
In the case that $|F|=O(1)$, i.e., the number of trees is assumed to be constant, it turns out that the assortment problem can also be solved efficiently, which we formalize in the proposition below.
Proposition 3.7
Assume $|F|=O(1)$ and $|\mathbf{leaves}(t)|$ is polynomial in $n$ for each tree $t\in F$. Then the assortment problem (2) can be solved optimally in a polynomial time.
Combined with Theorem 3.2, Proposition 3.7 shows that it is the number of trees $|F|$, rather than the structure of each tree (e.g., the number of leaves) that drives the complexity of the assortment problem (2) from being tractable to inapproximable (unless P = NP). Recall that in the reduction that proves Theorem 3.2, the number of leaf nodes in each tree is roughly $k^{2}/2$, which is $O(1)$ under the theorem’s assumption of constant $k$. Consequently, compared to Theorem 3.2, Proposition 3.7 suggests that even if the trees become significantly more complex – where the number of leaves in each tree increases from $O(1)$ to a polynomial in $n$ – the assortment optimization problem remains solvable in polynomial time, provided that the number of trees $|F|$ is $O(1)$.
4 Solution methodology based on Benders decomposition
While the formulations in Section 3 bring the assortment optimization problem under the decision forest choice model closer to being solvable in practice, the effectiveness of these formulations can be limited in large-scale problems. In particular, consider the case where there is a large number of trees in the decision forest model and each tree consists of a large number of splits and leaves. In this setting, both formulations – SplitMIO and ProductMIO– will have a large number of $y_{t,\ell}$ variables and a large number of constraints to link those variables with the $x_{i}$ variables, and may require significant computation time.
At the same time, SplitMIO and ProductMIO share a common problem structure. In particular, they have two sets of variables: the $\mathbf{x}$ variables, which determine the products that are to be included, and the $(\mathbf{y}_{t})_{t\in F}$ variables, which model the choice of each customer type. In addition, for any two trees $t,t^{\prime}$ such that $t\neq t^{\prime}$, the $\mathbf{y}_{t}$ variables and $\mathbf{y}_{t^{\prime}}$ variables do not appear together in any constraints. Thus, one can view each of our two formulations as a two-stage stochastic program, where each tree $t$ corresponds to a scenario; the variable $\mathbf{x}$ corresponds to the first-stage decision; and the variable $\mathbf{y}_{t}$ corresponds to the second-stage decision under scenario $t$, which is appropriately constrained by the first-stage decision $\mathbf{x}$.
Thus, one can apply Benders decomposition to solve the problem. At a high level, Benders decomposition involves using linear optimization duality to represent the optimal value of the second-stage problem for each tree $t$ as a piecewise-linear concave function of $\mathbf{x}$, and to eliminate the $(\mathbf{y}_{t})_{t\in F}$ variables. One can then re-write the optimization problem in epigraph form, resulting in an optimization problem in terms of the $\mathbf{x}$ variable and an auxiliary epigraph variable $\theta_{t}$ for each tree $t$, and a large family of constraints linking $\mathbf{x}$ and $\theta_{t}$ for each tree $t$. Although the constraints for each tree $t$ are too many to be enumerated, one can solve the problem through constraint generation.
The main message of this section is that the second-stage problem can be solved in a computationally efficient manner. In the remaining sections, we carefully analyze the second-stage problem for both formulations. For SplitMIO, we show that the second-stage problem can be solved by a greedy algorithm when $\mathbf{x}$ is fractional (Section 4.1). For ProductMIO, we show that the same greedy approach does not solve the second-stage problem in the fractional case (Section 4.2). For both formulations, when $\mathbf{x}$ is binary, we characterize the primal and dual solutions in closed form (Section 4.3). Lastly, in Section 4.4, we briefly describe our overall algorithmic approach to solving the assortment optimization problem, which involves solving the Benders reformulation of the relaxed problem, followed by the Benders reformulation of the integer problem.
4.1 Benders reformulation of the SplitMIO relaxation
The Benders reformulation of the LO relaxation of SplitMIO can be written as
$\underset{\mathbf{x},\boldsymbol{\theta}}{\text{maximize}}$ $\sum_{t\in F}\lambda_{t}\theta_{t}$ (9a)
subject to $\theta_{t}\leq G_{t}(\mathbf{x}),\quad\forall\ t\in F,$ (9b)
$\mathbf{x}\in[0,1]^{n},$ (9c)
where the function $G_{t}(\mathbf{x})$ is the optimal value of the following subproblem for tree $t$:
$G_{t}(\mathbf{x})=\quad$ $\underset{\mathbf{y}_{t}}{\text{maximize}}$ $\sum_{\ell\in\mathbf{leaves}(t)}r_{t,\ell}\cdot y_{t,\ell}$ (10a)
subject to $\sum_{\ell\in\mathbf{leaves}(t)}y_{t,\ell}=1,$ (10b)
$\sum_{\ell\in\mathbf{left}(s)}y_{t,\ell}\leq x_{v(t,s)},\quad\forall\ s\in\mathbf{splits}(t),$ (10c)
$\sum_{\ell\in\mathbf{right}(s)}y_{t,\ell}\leq 1-x_{v(t,s)},\quad\forall\ s\in\mathbf{splits}(t),$ (10d)
$y_{t,\ell}\geq 0,\quad\forall\ \ell\in\mathbf{leaves}(t).$ (10e)
We now present a greedy algorithm for solving problem (10) in Algorithm 1. This algorithm requires as input an ordering $\tau$ of the leaves in nondecreasing revenue. In particular, we require a bijection $\tau:\{1,\dots,|\mathbf{leaves}(t)|\}\to\mathbf{leaves}(t)$ such that $r_{t,\tau(1)}\geq r_{t,\tau(2)}\geq\dots\geq r_{t,\tau(|\mathbf{leaves}(t)|{\color[rgb]{0,0,0})}}$, i.e., an ordering of leaves in nondecreasing revenue. In addition, in the definition of Algorithm 1, we use $\mathbf{LS}(\ell)$ and $\mathbf{RS}(\ell)$ to denote the sets of left and right splits of $\ell$, respectively, which are defined as
$\mathbf{LS}(\ell)=\{s\in\mathbf{splits}(t)\mid\ell\in\mathbf{left}(s)\}\quad\text{and}\quad\mathbf{RS}(\ell)=\{s\in\mathbf{splits}(t)\mid\ell\in\mathbf{right}(s)\}.$
In words, $\mathbf{LS}(\ell)$ is the set of splits for which we must proceed to the left in order to be able to reach $\ell$, and $\mathbf{RS}(\ell)$ is the set of splits for which we must proceed to the right to reach $\ell$. A split $s\in\mathbf{LS}(\ell)$ if and only if $\ell\in\mathbf{left}(s)$, and similarly, $s\in\mathbf{RS}(\ell)$ if and only if $\ell\in\mathbf{right}(s)$.
Intuitively, this algorithm progresses through the leaves from highest to lowest revenue, and sets the $y_{t,\ell}$ variable of each leaf $\ell$ to the highest value it can be set to without violating the left and right split constraints (10c) and (10d) and without violating the constraint $\sum_{\ell\in\mathbf{leaves}(t)}y_{t,\ell}\leq 1$. At each iteration, the algorithm additionally keeps track of which constraint became tight through the event set $\mathcal{E}$. An $A_{s}$ event indicates that the left split constraint (10c) for split $s$ became tight; a $B_{s}$ event indicates that the right split constraint (10d) for split $s$ became tight; and a $C$ event indicates that the constraint $\sum_{\ell\in\mathbf{leaves}(t)}y_{t,\ell}\leq 1$ became tight. When a $C$ event is not triggered, Algorithm 1 looks for the split which has the least remaining capacity (line 15). In the case that the $\arg\min$ is not unique and there are two or more splits that are tied, we break ties by choosing the split $s$ with the lowest depth $d(s)$ (i.e., the split closest to the root node of the tree).
The function $f$ keeps track of which leaf $\ell$ was being checked when an $A_{s}$ / $B_{s}$ / $C$ event occurred. We note that both $\mathcal{E}$ and $f$ are not needed to find the primal solution, but they are essential to determining the dual solution in the dual procedure (Algorithm 2, which we will define shortly).
[…]
0:  Bijection $\tau:\{1,\dots,|\mathbf{leaves}(t)|\}\to\mathbf{leaves}(t)$ such that $r_{t,\tau(1)}\geq r_{t,\tau(2)}\geq\dots\geq r_{t,\tau(|\mathbf{leaves}(t)|)}$
1:  Initialize $y_{t,\ell}\leftarrow 0$ for each $\ell\in\mathbf{leaves}(t)$.
2:  for  $i=1,\dots,|\mathbf{leaves}(t)|$ do
3:     Set $q_{C}\leftarrow 1-\sum_{j=1}^{i-1}y_{t,\tau(j)}$.
4:     for $s\in\mathbf{LS}(\tau(i))$ do
5:        Set $q_{s}\leftarrow x_{v(t,s)}-{\color[rgb]{0,0,0}\sum_{1\leq j\leq i-1:\tau(j)\in\mathbf{left}(s)}y_{t,\tau(j)}}$
6:     for $s\in\mathbf{RS}(\tau(i))$ do
7:        Set $q_{s}\leftarrow 1-x_{v(t,s)}-{\color[rgb]{0,0,0}\sum_{1\leq j\leq i-1:\tau(j)\in\mathbf{right}(s)}y_{t,\tau(j)}}$
8:     Set $q_{A,B}\leftarrow\min_{s\in\mathbf{LS}(\tau(i))\cup\mathbf{RS}(\tau(i))}q_{s}$
9:     Set $q^{*}\leftarrow\min\{q_{C},q_{A,B}\}$
10:     Set $y_{t,\tau(i)}\leftarrow q^{*}$
11:     if  $q^{*}=q_{C}$  then
12:        Set $\mathcal{E}\leftarrow\mathcal{E}\cup\{C\}$.
13:        Set $f(C)=\tau(i)$.
14:     else
15:        Set $s^{*}\leftarrow\arg\min_{s\in\mathbf{LS}(\tau(i))\cup\mathbf{RS}(\tau(i))}q_{s}$
16:        if  $s^{*}\in\mathbf{LS}(\tau(i))$  then
17:           Set $e=A_{{\color[rgb]{0,0,0}s^{*}}}$
18:        else
19:           Set $e=B_{{\color[rgb]{0,0,0}s^{*}}}$
20:        if $e\notin\mathcal{E}$ then
21:           Set $\mathcal{E}\leftarrow\mathcal{E}\cup\{e\}$.
22:           Set $f(e)=\tau(i)$.
Algorithm 1 Primal greedy algorithm for SplitMIO.
It turns out that Algorithm 1 returns a feasible solution that is an extreme point of the polyhedron defined in problem (10). We state the result formally as the following theorem.
Theorem 4.1
Fix $t\in F$. Let $\mathbf{y}_{t}$ be a solution to problem (10) produced by Algorithm 1. Then: (a) $\mathbf{y}_{t}$ is a feasible solution to problem (10); and (b) $\mathbf{y}_{t}$ is an extreme point of the feasible region of problem (10).
Theorem 4.1 implies that problem (10) is feasible, and since the problem is bounded, it has a finite optimal value. By strong duality, the optimal value of problem (10) is equal to the optimal value of its dual:
$\underset{\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t}}{\text{minimize}}$ $\sum_{s\in\mathbf{splits}(t)}x_{v(t,s)}\cdot\alpha_{t,s}+\sum_{s\in\mathbf{splits}(t)}(1-x_{v(t,s)})\beta_{t,s}+\gamma_{t}$ (11a)
subject to $\sum_{s:\ell\in\mathbf{left}(s)}\alpha_{t,s}+\sum_{s:\ell\in\mathbf{right}(s)}\beta_{t,s}+\gamma_{t}\geq r_{t,\ell},\quad\forall\ \ell\in\mathbf{leaves}(t),$ (11b)
$\alpha_{t,s}\geq 0,\quad\forall\ s\in\mathbf{splits}(t),$ (11c)
$\beta_{t,s}\geq 0,\quad\forall\ s\in\mathbf{splits}(t).$ (11d)
Letting $\mathcal{D}_{t,\textsc{SplitMIO}}$ denote the set of feasible solutions to the dual subproblem (11), we can formulate the master problem (9) as
$\underset{\mathbf{x},\boldsymbol{\theta}}{\text{maximize}}$ $\sum_{t\in F}\lambda_{t}\theta_{t}$ (12a)
subject to $\theta_{t}\leq\sum_{s\in\mathbf{splits}(t)}x_{v(t,s)}\cdot\alpha_{t,s}+\sum_{s\in\mathbf{splits}(t)}(1-x_{v(t,s)})\beta_{t,s}+\gamma_{t},$
$\quad\forall\ (\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})\in\mathcal{D}_{t,\textsc{SplitMIO}},$ (12b)
$\mathbf{x}\in[0,1]^{n}.$ (12c)
The value of this formulation, relative to the original formulation, is that we have replaced the $(\mathbf{y}_{t})_{t\in F}$ variables and the constraints that link them to the $\mathbf{x}$ variables, with a large family of constraints in terms of $\mathbf{x}$. Although this new formulation is still challenging, the advantage of this formulation is that it is suited to constraint generation.
The constraint generation approach to solving problem (12) involves starting the problem with no constraints and then, for each $t\in F$, checking whether constraint (12b) is violated. If constraint (12b) is not violated for any $t\in F$, then we conclude that the current solution $\mathbf{x}$ is optimal. Otherwise, for any $t\in F$ such that constraint (12b) is violated, we add the constraint corresponding to the $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ solution at which the violation occurred, and solve the problem again to obtain a new $\mathbf{x}$. The procedure then repeats at the new $\mathbf{x}$ solution until no more violated constraints have been found.
The crucial step to solving this problem is being able to solve the dual subproblem (11); that is, for a fixed $t\in F$, either asserting that the current solution $\mathbf{x}$ satisfies constraint (12b) for all $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})\in\mathcal{D}_{t,\textsc{SplitMIO}}$ or identifying a $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ at which constraint (12b) is violated. This amounts to solving the dual subproblem (11) and comparing its objective value to $\theta_{t}$.
Fortunately, it turns out that we can solve the dual subproblem (11) using a specialized algorithm, in the same way that we can solve the primal subproblem (10) using Algorithm 1. Algorithm 2 uses auxiliary information obtained during the execution of Algorithm 1. In the definition of Algorithm 2, we use $d(s)$ to denote the depth of an arbitrary split, where the root split corresponds to a depth of 1, and $d_{\max}=\max_{s\in\mathbf{splits}(t)}d(s)$ is the depth of the deepest split in the tree. In addition, we use $\mathbf{splits}(t,d)=\{s\in\mathbf{splits}(t)\mid d(s)=d\}$ to denote the set of all splits at a particular depth $d$. Algorithm 2 proceeds sequentially from the root to the deepest splits, updating the dual variables $\alpha_{t,s}$ and $\beta_{t,s}$ for each split $s$ based on the tree topology and the auxiliary information $\mathcal{E}$ and $f$ obtained from Algorithm 1. When $A_{s}\in\mathcal{E}$, constraint (10c) in the primal problem is tight, whereas constraint (10d) is not. This implies that the dual variable $\beta_{t,s}$, associated with the latter constraint, must remain zero. The other dual variable, $\alpha_{t,s}$, is then updated according to constraint (11b) in the dual problem. Specifically, the auxiliary function $f$ in Algorithm 1 tracks the leaf node $\ell=f(A_{s})$ with nonzero $y_{t,\ell}$ when $A_{s}$ occurs. Since the corresponding constraint for $\ell$ in constraint (11b) is tight, it is used to calibrate $\alpha_{t,s}$. A symmetric argument applies when $B_{s}\in\mathcal{E}$ instead. We provide a worked example of the execution of both Algorithms 1 and 2 in Appendix 9.
[…]
1:  Initialize $\alpha_{t,s}\leftarrow 0,\beta_{t,s}\leftarrow 0$ for all $s\in\mathbf{splits}(t)$, $\gamma_{t}\leftarrow 0$
2:  Set $\gamma\leftarrow r_{f(C)}$
3:  for  $d=1,\dots,d_{\max}$  do
4:     for  $s\in\mathbf{splits}(t,d)$  do
5:        if  $A_{s}\in\mathcal{E}$ then
6:           Set $\alpha_{t,s}\leftarrow r_{t,f(A_{s})}-\gamma_{t}-\sum_{\begin{subarray}{c}s^{\prime}\in\mathbf{LS}(f(A_{s})):A_{s^{\prime}}\in\mathcal{E},d(s^{\prime})<d\end{subarray}}\alpha_{t,s^{\prime}}-\sum_{\begin{subarray}{c}s^{\prime}\in\mathbf{RS}(f(A_{s})):B_{s^{\prime}}\in\mathcal{E},d(s^{\prime})<d\end{subarray}}\beta_{t,s^{\prime}}$
7:        if  $B_{s}\in\mathcal{E}$ then
8:           Set $\beta_{t,s}\leftarrow r_{t,f(B_{s})}-\gamma_{t}-\sum_{\begin{subarray}{c}s^{\prime}\in\mathbf{LS}(f(A_{s})):A_{s^{\prime}}\in\mathcal{E},d(s^{\prime})<d\end{subarray}}\alpha_{t,s^{\prime}}-\sum_{\begin{subarray}{c}s^{\prime}\in\mathbf{RS}(f(A_{s})):B_{s^{\prime}}\in\mathcal{E},d(s^{\prime})<d\end{subarray}}\beta_{t,s^{\prime}}$
Algorithm 2 Dual greedy algorithm for SplitMIO.
We can show that the dual solution produced by Algorithm 2 is a feasible extreme point solution of the dual subproblem (11).
Theorem 4.2
Fix $t\in F$. Let $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ be a solution to problem (11) produced by Algorithm 2. Then: (a) $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ is a feasible solution to problem (11); and (b) $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ is an extreme point of the feasible region of problem (11).
Lastly, and most importantly, we show that the solutions produced by Algorithms 1 and 2 are optimal for their respective problems. Thus, Algorithm 2 is a valid procedure for identifying values of $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ at which constraint (12b) is violated.
Theorem 4.3
Fix $t\in F$. Let $\mathbf{y}_{t}$ be a solution to problem (10) produced by Algorithm 1 and $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ be a solution to problem (11) produced by Algorithm 2. Then: (a) $\mathbf{y}_{t}$ is an optimal solution to problem (10); and (b) $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ is an optimal solution to problem (11).
The proof of this result follows by verifying that the two solutions satisfy complementary slackness.
Before continuing, we pause to make two remarks. First, Algorithms 1 and 2 can be viewed as the generalization of the algorithms arising in the Benders decomposition approach to the ranking-based assortment optimization problem (Bertsimas and Mišić 2019). The results of that paper show that the primal subproblem of the MIO formulation in Bertsimas and Mišić (2019) can be solved via a greedy algorithm (analogous to Algorithm 1) and the dual subproblem can be solved via an algorithm that uses information from the primal algorithm (analogous to Algorithm 2). This generalization is not straightforward. The main challenge in this generalization is designing the sequence of updates in the greedy algorithm according to the tree topology. For example, in Algorithm 1, one considers all left/right splits and the $y$ values of their left/right leaves when constructing the lowest upper bound of $y_{\ell}$ for each leaf node $\ell$. Also, as shown in Algorithm 2, the dual variables $\alpha_{t,s}$ and $\beta_{t,s}$ have to be updated according to the tree topology and the events $A_{s^{\prime}}$ and $B_{s^{\prime}}$ of the split $s^{\prime}$ with smaller depth. In contrast, in the ranking-based assortment optimization problem, one only needs to calculate the “capacities” (the $q_{s}$ values in Algorithm 1) by subtracting the $y$ values of the preceding products in the ranking, which does not inherit any topological structure. For these reasons, the primal and dual Benders subproblems for the decision forest assortment problem are much more challenging than those of the ranking-based assortment problem.
Second, while the dual problem (11) can be solved directly using commercial software such as Gurobi (Gurobi Optimization, LLC 2024), the proposed greedy algorithm is generally more efficient. In Appendix 11.3, we numerically demonstrate its efficiency, showing that when solving the relaxed SplitMIO problem, the greedy algorithm can reduce runtime by up to 90% compared to the approach where each constraint is added via solving the dual subproblem using Gurobi.
4.2 Benders reformulation of the ProductMIO relaxation
We also consider a Benders reformulation of the linear relaxation of ProductMIO. The Benders master problem is given by formulation (9) where the function $G_{t}(\mathbf{x})$ is defined as the optimal value of the ProductMIO subproblem for tree $t$. To aid in the definition of the subproblem, let $P(t)$ denote the set of products that appear in the splits of tree $t$:
$P(t)=\{i\in\mathcal{N}\mid i=v(t,s)\ \text{for some}\ s\in\mathbf{splits}(t)\}.$
With a slight abuse of notation, let $\mathbf{left}(i)$ denote the set of leaves for which product $i$ must be included in the assortment for those leaves to be reached, and similarly, let $\mathbf{right}(i)$ denote the set of leaves for which product $i$ must be excluded from the assortment for those leaves to be reached; formally,
$\mathbf{left}(i)=\bigcup_{\begin{subarray}{c}s\in\mathbf{splits}(t):v(t,s)=i\end{subarray}}\mathbf{left}(s),\qquad\mathbf{right}(i)=\bigcup_{\begin{subarray}{c}s\in\mathbf{splits}(t):v(t,s)=i\end{subarray}}\mathbf{right}(s).$
With these definitions, we can write down the ProductMIO subproblem as follows:
$G_{t}(\mathbf{x})=\quad$ $\underset{\mathbf{y}_{t}}{\text{maximize}}$ $\sum_{\ell\in\mathbf{leaves}(t)}r_{t,\ell}\cdot y_{t,\ell}$ (13a)
subject to $\sum_{\ell\in\mathbf{leaves}(t)}y_{t,\ell}=1,$ (13b)
$\sum_{\ell\in\mathbf{left}(i)}y_{t,\ell}\leq x_{i},\quad\forall\ i\in P(t),$ (13c)
$\sum_{\ell\in\mathbf{right}(i)}y_{t,\ell}\leq 1-x_{i},\quad\forall\ i\in P(t),$ (13d)
$y_{t,\ell}\geq 0,\quad\forall\ \ell\in\mathbf{leaves}(t).$ (13e)
In the same way as SplitMIO, one can consider solving problem (13) using a greedy approach, where one iterates through the leaves from highest to lowest revenue, and sets each leaf’s $y_{t,\ell}$ variable to the highest possible value without violating any of the constraints. Unlike SplitMIO, it unfortunately turns out that this greedy approach is not always optimal. We formalize this claim as a proposition in Section 10 of the appendix, supported by a counterexample.
4.3 Bender Cuts for Integer Master Solutions
We further propose closed form expressions for the structure of the optimal primal and dual Benders subproblem solutions for integer solutions $\mathbf{x}$.
4.3.1 SplitMIO
Our results in Section 4.1 for obtaining primal and dual solutions for the subproblem of SplitMIO apply for any $\mathbf{x}\in[0,1]^{n}$; in particular, they apply for fractional choices of $\mathbf{x}$, thus allowing us to solve the Benders reformulation of the relaxation of SplitMIO.
In the special case that $\mathbf{x}$ is a candidate integer solution of SplitMIO, we can find optimal solutions to the primal and dual subproblems of SplitMIO in closed form.
Theorem 4.4
Fix $t\in F$, and let $\mathbf{x}\in\{0,1\}^{n}$. Define the primal subproblem solution $\mathbf{y}_{t}$ as
$y_{t,\ell}=\left\{\begin{array}[]{ll}1&\text{if}\ \ell=\ell^{*},\\ 0&\text{if}\ \ell\neq\ell^{*},\end{array}\right.$
where $\ell^{*}$ denotes the leaf that the assortment encoded by $\mathbf{x}$ is mapped to. Define the dual subproblem solution $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ as
$\alpha_{t,s}$ =$=\left\{\begin{array}[]{ll}\max\{0,\max_{\ell\in\mathbf{left}(s)}r_{t,\ell}-r_{t,\ell^{*}}\}&\text{if}\ s\in\mathbf{RS}(\ell^{*}),\\ 0&\text{otherwise},\end{array}\right.$
$\beta_{t,s}$ =$=\left\{\begin{array}[]{ll}\max\{0,\max_{\ell\in\mathbf{right}(s)}r_{t,\ell}-r_{t,\ell^{*}}\}&\text{if}\ s\in\mathbf{LS}(\ell^{*}),\\ 0&\text{otherwise},\end{array}\right.$
$\gamma_{t}$ $=r_{t,\ell^{*}}.$
Then: (a) $\mathbf{y}_{t}$ is a feasible solution to problem (10); (b) $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ is a feasible solution to problem (11); and (c) $\mathbf{y}_{t}$ and $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ are optimal for problems (10) and (11), respectively.
The significance of Theorem 4.4 is that it provides a simpler means to checking for violated constraints when $\mathbf{x}$ is binary than applying Algorithms 1 and 2. In particular, for the integer version of SplitMIO, a similar derivation as in Section 4.1 leads us to the following Benders reformulation of the integer problem for the SplitMIO formulation:
$\underset{\mathbf{x},\boldsymbol{\theta}}{\text{maximize}}$ $\sum_{t\in F}\lambda_{t}\theta_{t}$ (14a)
subject to $\theta_{t}\leq\sum_{s\in\mathbf{splits}(t)}x_{v(t,s)}\cdot\alpha_{t,s}+\sum_{s\in\mathbf{splits}(t)}(1-x_{v(t,s)})\beta_{t,s}+\gamma_{t},$
$\quad\forall\ (\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})\in\mathcal{D}_{t,\textsc{SplitMIO}},$ (14b)
$\mathbf{x}\in\{0,1\}^{n}.$ (14c)
To check whether constraint (14b) is violated for a particular $\mathbf{x}$ and a tree $t$, we can simply use Theorem 4.4 to determine the optimal value of the subproblem, and compare it against $\theta_{t}$; if the constraint corresponding to the dual solution of Theorem 4.4 is violated, we add that constraint to the problem. In our implementation of Benders decomposition, we embed the constraint generation process for the integer problem (14) within the branch-and-bound tree, using a technique referred to as lazy constraint generation; we discuss this more in Section 4.4.
4.3.2 ProductMIO
We also consider ProductMIO. We begin by writing down the dual of the subproblem, for which we need to define several additional sets. We let $\mathbf{LP}(\ell)$ denote the set of “left products” of leaf $\ell$ (those products that must be included in the assortment for leaf $\ell$ to be reached), and let $\mathbf{RP}(\ell)$ denote the set of “right products” of leaf $\ell$ (those products that must be excluded from the assortment for leaf $\ell$ to be reached). Note that $\ell\in\mathbf{left}(i)$ if and only if $i\in\mathbf{LP}(\ell)$, and similarly $\ell\in\mathbf{right}(i)$ if and only if $i\in\mathbf{RP}(\ell)$.
With these definitions, the dual of the primal subproblem (13) is
$\underset{\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t}}{\text{minimize}}$ $\sum_{i\in P(t)}\alpha_{t,i}x_{i}+\sum_{i\in P(t)}\beta_{t,i}(1-x_{i})+\gamma_{t}$ (15a)
subject to $\sum_{i\in\mathbf{LP}(\ell)}\alpha_{t,i}+\sum_{i\in\mathbf{RP}(\ell)}\beta_{t,i}+\gamma_{t}\geq r_{\ell},\quad\forall\ \ell\in\mathbf{leaves}(t),$ (15b)
$\alpha_{t,i}\geq 0,\quad\forall\ i\in P(t),$ (15c)
$\beta_{t,i}\geq 0,\quad\forall\ i\in P(t).$ (15d)
In the case that $\mathbf{x}$ is integer, we can obtain optimal solutions to the primal subproblem (13) and its dual (15) in closed form.
Theorem 4.5
Fix $t\in F$ and let $\mathbf{x}\in\{0,1\}^{n}$. Let $\mathbf{y}_{t}$ be defined as in Theorem 4.4 and let $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ be defined as
$\alpha_{t,i}$ =$=\left\{\begin{array}[]{ll}\max\{0,\max_{\ell\in\mathbf{left}(i)}r_{t,\ell}-r_{t,\ell^{*}}\},&\text{if}\ i\in\mathbf{RP}(\ell^{*}),\\ 0&\text{otherwise},\end{array}\right.$ (18)
$\beta_{t,i}$ =$=\left\{\begin{array}[]{ll}\max\{0,\max_{\ell\in\mathbf{right}(i)}r_{t,\ell}-r_{t,\ell^{*}}\},&\text{if}\ i\in\mathbf{LP}(\ell^{*}),\\ 0&\text{otherwise},\end{array}\right.$ (21)
$\gamma_{t}$ $=r_{t,\ell^{*}}.$ (22)
Then: (a) $\mathbf{y}_{t}$ is a feasible solution for problem (13); (b) $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ is a feasible solution for problem (15); and (c) $\mathbf{y}_{t}$ and $(\boldsymbol{\alpha}_{t},\boldsymbol{\beta}_{t},\gamma_{t})$ are optimal for problems (13) and (15), respectively.
4.4 Overall Benders algorithm
We conclude Section 4 by summarizing how the results are used. In our overall algorithmic approach below, we first focus on SplitMIO, as the subproblem can be solved for the formulation when $\mathbf{x}$ is either fractional or binary.
  1. 1.
    Relaxation phase. We first solve the relaxed problem (12) using ordinary constraint generation. Given a solution $\mathbf{x}\in[0,1]^{n}$, we generate Benders cuts by running the primal-dual procedure ( Algorithm 1 followed by Algorithm 2).
  2. 2.
    Integer phase. In the integer phase, we add all of the Benders cuts generated in the relaxation phase to the integer version of problem (12). We then solve the problem as an integer optimization problem, where we generate Benders cuts for integer solutions using the closed form expressions in Theorem 4.4. We add these cuts using lazy constraint generation. That is, we solve the master problem using a single branch-and-bound tree, and we check whether the main constraint (12b) of the Benders formulation is violated at every integer solution generated in the branch-and-bound tree.
For ProductMIO, as the subproblem can only be solved when $\mathbf{x}$ is binary, its overall Benders algorithm directly starts with the Integer Phase.
We conclude this section by highlighting the advantages of the SplitMIO formulation in the large-scale regime. At first glance, following Section 3, ProductMIO may appear preferable to SplitMIO since it is a tighter formulation (Proposition 3.5) with fewer constraints. However, in large-scale settings, the proposed primal-dual greedy algorithm enables significantly faster computation of the linear relaxation for SplitMIO compared to ProductMIO. As a result, as outlined above, SplitMIO can benefit from a more efficient warm start during the relaxation phase. Additionally, the objective value of the linear relaxation provides a valid upper bound for the assortment problem (2). This upper bound is particularly valuable in large-scale instances where solving SplitMIO or ProductMIO exactly may be impractical. A strong upper bound helps reduce the number of branches explored in the branch-and-bound procedure and serves as a benchmark for assessing the suboptimality of a given assortment solution. In Appendix 11.3, we demonstrate that, with the greedy algorithm, solving the relaxed SplitMIO problem can be up to 95% faster than solving the relaxed ProductMIO while providing comparable upper bounds.
5 Numerical Experiments with Synthetic Data
In this section, we examine the formulation strength of SplitMIO and ProductMIO (Section 5.2) and demonstrate the scalability of the Benders decomposition approach(Section 5.3). We focus on synthetically generated instances so that we can scale the problem size. In Appendix 11, we report our implementation details and include additional numerical results.
5.1 Background
To test our method, we generate three different families of synthetic decision forest instances, which differ in the topology of the trees and the products that appear in the splits:
  1. 1.
    T1 instances. A T1 forest consists of balanced trees of depth $d$ (i.e., trees where all leaves are at depth $d+1$). For each tree, we sample $d$ products $i_{1},\dots,i_{d}$ uniformly without replacement from $\mathcal{N}$. Then, for every depth $d^{\prime}\in\{1,\dots,d\}$, we set the split product $v(t,s)$ as $v(t,s)=i_{d^{\prime}}$ for every split $s$ that is at depth $d^{\prime}$.
  2. 2.
    T2 instances. A T2 forest consists of balanced trees of depth $d$. For each tree, we set the split products at each split iteratively, starting at the root, in the following manner: (a) Initialize $d^{\prime}=1$; (b) For all splits $s$ at depth $d^{\prime}$, set $v(s,t)=i_{s}$ where $i_{s}$ is drawn uniformly at random from the set $\mathcal{N}\setminus\cup_{s^{\prime}\in A(s)}\{v(t,s^{\prime})\}$, where $A(s)$ is the set of ancestor splits to split $s$ (i.e., all splits appearing on the path from the root node to split $s$); (c) Increment $d^{\prime}\leftarrow d^{\prime}+1$; and (d) If $d^{\prime}>d$, stop; otherwise, return to Step (b).
  3. 3.
    T3 instances. A T3 forest consists of unbalanced trees with $L$ leaves. Each tree is generated according to the following iterative procedure: (a) Initialize $t$ to a tree consisting of a single leaf; (b) Select a leaf $\ell$ uniformly at random from $\mathbf{leaves}(t)$, and replace it with a split $s$ and two child leaves $\ell_{1},\ell_{2}$. For split $s$, set $v(s,t)=i_{s}$ where $i_{s}$ is drawn uniformly at random from $\mathcal{N}\setminus\cup_{s^{\prime}\in A(s)}\{v(t,s^{\prime})\}$; (c) If $|\mathbf{leaves}(t)|=L$, terminate; otherwise, return to Step (b).
For all three types of forests, we generate the purchase decision $c(t,\ell)$ for each leaf $\ell$ in each tree $t$ in the following way: for each leaf $\ell$, we uniformly at random choose a product $i\in\cup_{s\in\mathbf{LS}(\ell)}\{v(t,s)\}\cup\{0\}$. In words, the purchase decision is chosen to be consistent with the products that are known to be in the assortment if leaf $\ell$ is reached. Figure 2 shows an example of each type of tree (T1, T2, and T3). Given a forest of any of the three types above, we generate the customer type probability vector $\boldsymbol{\lambda}=(\lambda_{t})_{t\in F}$ by drawing it uniformly from the $(|F|-1)-$dimensional unit simplex.
(a) T1 tree ($d=3$).
(b) T2 tree ($d=3$).
(c) T3 tree ($L=8$).
Figure 2: Examples of T1, T2 and T3 trees.
In our experiments, we fix the number of products $n=100$ and vary the number of trees $|F|\in\{50,100,200,500\}$, and the number of leaves $|\mathbf{leaves}(t)|\in\{8,16,32,64\}$. (Note that the chosen values for $|\mathbf{leaves}(t)|$ correspond to depths of $\{3,4,5,6\}$ for the T1 and T2 instances.) For each combination of $n$, $|F|$ and $|\mathbf{leaves}(t)|$ and each type of instance (T1, T2 and T3) we randomly generate 20 problem instances, where a problem instance consists of a decision forest model and the product marginal revenues $\rho_{1},\dots,\rho_{n}$. For each instance, the decision forest model is generated according to the process described above and the product revenues are sampled uniformly with replacement from the set $\{1,\dots,100\}$.
5.2 Experiment #1: Formulation Strength
Our first experiment is to simply understand how the two formulations – SplitMIO and ProductMIO– compare in terms of formulation strength. Recall from Proposition 3.5 that ProductMIO is at least as strong as SplitMIO. For a given instance and a given formulation $\mathcal{M}$ (one of SplitMIO and ProductMIO), we define the integrality gap $G^{\text{int}}_{\mathcal{M}}\equiv 100\%\times\left({Z_{\mathcal{M}}-Z^{*}}\right)/{Z^{*}}$, where $Z^{*}$ is the optimal objective value of the integer problem and $Z_{\mathcal{M}}$ is the optimal objective of the LO relaxation. We consider the T1, T2 and T3 instances with $n=100$, $|F|\in\{50,100,200,500\}$ and $|\mathbf{leaves}(t)|=8$. We restrict our focus to instances with $n=100$ and $|\mathbf{leaves}(t)|=8$, as the optimal value $Z^{*}$ of the integer problem could be computed within one hour for these instances.
Type T1 Type T2 Type T3
$|F|$ $G^{\text{int}}_{\textsc{SplitMIO}}$ $G^{\text{int}}_{\textsc{ProductMIO}}$ $|F|$ $G^{\text{int}}_{\textsc{SplitMIO}}$ $G^{\text{int}}_{\textsc{ProductMIO}}$ $|F|$ $G^{\text{int}}_{\textsc{SplitMIO}}$ $G^{\text{int}}_{\textsc{ProductMIO}}$
50 0.9 0.0 50 0.2 0.2 50 0.2 0.2
100 2.5 0.1 100 1.0 1.0 100 0.5 0.5
200 5.6 0.2 200 5.4 5.3 200 4.1 3.9
500 15.8 3.3 500 16.7 16.4 500 14.2 14.0
Table 1: Average integrality gap of SplitMIO and ProductMIO for T1, T2 and T3 instances.
Table 1 displays the average integrality gap of each of the two formulations for each combination of $n$ and $|F|$ and each instance type. From this table, we can see that the integrality gap of both SplitMIO and ProductMIO is in general about 0 to 17%. Note that the difference between ProductMIO and SplitMIO is most pronounced for the T1 instances, as the decision forests in these instances exhibit the highest degree of repetition of products within the splits of a tree. In contrast, the difference is much smaller for the T2 and T3 instances, where there is less repetition of products within the splits of the tree (as the trees are not forced to have the same product appear on all of the splits at a particular depth). Note that these results align with our theoretical analysis of ProductMIO and SplitMIO (specifically Proposition 3.5). Note also that when a tree does not have any repeated products in its splits, the corresponding constraints of ProductMIO and SplitMIO are the same, and ProductMIO does not confer an advantage over SplitMIO. Thus, as one constructs trees with less and less repetition in the splits, the improvement in the integrality gap of ProductMIO over SplitMIO should become smaller, with the two formulations having the same gap in the most extreme case where no product is repeated.
We also test the tractability of SplitMIO and ProductMIO when they are solved as integer programs (i.e., not as relaxation). We present the results in Section 11.1.
5.3 Experiment #2: Benders Decomposition for Large-Scale Problems
In this experiment, we report on the performance of our Benders decomposition approach for solving large scale instances of SplitMIO. We focus on the SplitMIO formulation, as we are able to efficiently generate Benders cuts for both fractional and integral values of $\mathbf{x}$.
We deviate from our previous experiments by generating a collection of T3 instances with $n\in\{200,500,1000,2000,3000\}$, $|F|=500$ trees and $|\mathbf{leaves}(t)|=512$ leaves. As before, the marginal revenue $\rho_{i}$ of each product $i$ is chosen uniformly at random from $\{1,\dots,100\}$. For each value of $n$, we generate 5 instances. For each instance, we solve the SplitMIO problem subject to the constraint $\sum_{i=1}^{n}x_{i}=b$, where $b$ is set as $b=\mu n$ and we vary $\mu\in\{0.02,0.04,0.06,0.08,0.10,0.12\}$.
We compare three different methods: the two-phase Benders method described in Section 4.4, using the SplitMIO cut results (Section 4.1 and Section 4.3); the divide-and-conquer (D&C) heuristic; and the direct solution approach, where we attempt to directly solve the full SplitMIO formulation using Gurobi (Gurobi Optimization, LLC 2024). The D&C heuristic is a type of local search heuristic proposed in the product line design literature (see Green and Krieger 1993; see also Belloni et al. 2008). In this heuristic, one iterates through the $b$ products currently in the assortment, and replaces a single product with the product outside of the assortment that leads to the highest improvement in the expected revenue; this process repeats until the assortment can no longer be improved. We choose the initial assortment uniformly at random from the collection of assortments of size $b$. For each instance, we repeat the D&C heuristic 10 times, and retain the best solution. We do not impose a time limit on the D&C heuristic. For the Benders approach, we impose a time limit of one hour on the LO phase, and impose a time limit of one hour on the integer phase. For the direct solution approach, we impose a time limit of two hours, in order to be comparable to the Benders approach.
Table 2 reports the performance of the three methods – the Benders approach, the D&C heuristic and direct solution of SplitMIO– across all combinations of $n$ and $\mu$. We consider several metrics. The metric $G_{\mathcal{M}}$ is defined as $G_{\mathcal{M}}=(Z^{\prime}-Z_{\mathcal{M}})/Z^{\prime}\times 100\%$, i.e., it is the optimality gap of the solution obtained by method $\mathcal{M}$ – either the Benders approach, the direct approach or the D&C heuristic – relative to the objective value of the best solution obtained out of the three methods, which is indicated by $Z^{\prime}$. Lower values of $G_{\mathcal{M}}$ indicate that the approach tends to deliver solutions that are close to being the best out of the three methods, and a value of 0% implies that the solution obtained by the approach is the best (or tied for the best) out of the three methods. The metric $T_{\mathcal{M}}$ indicates the solution time required for approach $\mathcal{M}$ in seconds. Finally, for the Benders and direct approaches, we compute the optimality gap $O_{\mathcal{M}}$, which is defined as $O_{\mathcal{M}}=(Z_{\text{UB},\mathcal{M}}-Z_{\text{LB},\mathcal{M}})/Z_{\text{UB},\mathcal{M}}\times 100\%$, where $Z_{\text{UB},\mathcal{M}}$ and $Z_{\text{LB},\mathcal{M}}$ are the best upper and lower bounds, respectively, obtained from either the Benders or direct approach after the computation time limit is exhausted. The value reported of each metric is the average over the five replications corresponding to the particular $(n,\mu)$ combination.
$n$ $\mu$ $b$ $G_{\text{Benders}}$ $G_{\text{D\&C}}$ $G_{\text{Direct}}$ $T_{\text{Benders}}$ $T_{\text{D\&C}}$ $T_{\text{Direct}}$ $O_{\text{Benders}}$ $O_{\text{Direct}}$
200 0.02 4 0.00 0.00 0.00 14.54 2.86 2060.81 0.00 0.00
200 0.04 8 0.00 0.04 0.61 187.96 4.46 7015.65 0.00 7.72
200 0.06 12 0.11 0.00 0.77 3615.82 7.34 7200.25 8.23 16.71
200 0.08 16 0.70 0.00 2.86 3612.43 9.97 7200.47 17.38 23.39
200 0.10 20 0.47 0.01 6.94 3613.19 15.07 7200.16 22.13 30.38
200 0.12 24 0.67 0.12 7.68 3614.96 18.93 7200.25 27.27 34.03
500 0.02 10 0.00 0.19 0.00 16.71 11.24 184.29 0.00 0.00
500 0.04 20 0.00 0.20 0.08 1640.78 28.93 7200.21 0.48 3.06
500 0.06 30 0.02 0.68 3.57 3630.19 65.75 7200.26 8.10 11.89
500 0.08 40 0.00 0.48 1.05 3623.51 115.36 7200.57 14.26 14.82
500 0.10 50 0.00 1.49 6.88 3625.89 177.26 7200.18 18.26 23.74
500 0.12 60 0.69 0.95 4.11 3631.14 223.61 7200.38 22.94 25.30
1000 0.02 20 0.00 0.29 0.00 18.83 47.62 156.49 0.00 0.00
1000 0.04 40 0.00 1.65 2.76 2823.19 183.33 7200.18 1.05 4.10
1000 0.06 60 0.00 2.48 6.61 3671.12 397.35 7200.18 6.16 12.53
1000 0.08 80 0.00 2.92 7.86 3662.76 687.02 7200.44 10.07 17.39
1000 0.10 100 0.00 2.31 8.65 3670.95 1052.76 7200.17 13.60 20.99
1000 0.12 120 0.00 2.13 5.92 3696.24 1552.76 7200.22 15.77 20.73
2000 0.02 40 0.00 1.36 0.00 14.55 328.25 70.02 0.00 0.00
2000 0.04 80 0.00 3.49 0.00 1774.85 1259.21 1057.28 0.21 0.00
2000 0.06 120 0.00 4.26 21.38 3774.67 2578.07 7200.16 2.42 23.29
2000 0.08 160 0.00 4.63 100.00 3806.42 4431.39 7200.28 5.21 100.00
2000 0.10 200 0.00 5.23 100.00 3925.28 6435.67 7200.17 7.16 100.00
2000 0.12 240 0.74 0.94 100.00 4319.62 10949.83 7200.16 11.91 100.00
3000 0.02 60 0.00 1.97 0.00 16.16 923.12 32.40 0.00 0.00
3000 0.04 120 0.00 4.03 0.00 1883.80 3365.35 1541.40 0.08 0.00
3000 0.06 180 0.00 5.26 0.04 4144.41 7620.89 7200.16 1.81 1.80
3000 0.08 240 0.00 4.55 99.98 4554.74 13624.07 7209.41 4.23 99.99
3000 0.10 300 0.00 4.04 99.98 5013.31 31137.18 7200.38 6.17 99.98
3000 0.12 360 0.32 1.12 99.98 6999.60 43173.66 7216.75 9.83 99.99
Table 2: Comparison of the Benders decomposition approach, the D&C heuristic and direct solution of SplitMIO in terms of solution quality, computation time and optimality gap.
Comparing the performance of the Benders approach with the D&C heuristic, we can see that in general, the Benders approach is able to find better solutions than the D&C heuristic. In particular, for larger instances, $G_{\text{Benders}}$ is lower than $G_{\text{D\&C}}$ (for example, with $n=2000$, $\mu=0.08$, the Benders solution is in general the best one, and the D&C solution has an expected revenue that is 4.6% worse). In addition, from a computation time standpoint, the Benders approach compares quite favorably to the D&C heuristic. While the D&C heuristic is faster for small problems with low $n$ and/or low $\mu$, it can require a significant amount of time for $n=2000$ or $n=3000$. In addition to this comparison against the D&C heuristic, in Appendix 11.2, we also provide a comparison of the MIO solutions for the smaller T1, T2 and T3 instances used in the previous two sections against three other heuristic solutions; in those instances, we similarly find that solutions obtained from our MIO formulations can be significantly better than heuristic solutions.
Comparing the performance of the Benders approach with the direct solution approach, our results indicate two types of behavior. The first type of behavior corresponds to “easy” instances. These are instances with $\mu\in\{0.02,0.04\}$ for which it is sometimes possible to directly solve SplitMIO to optimality within the two hour time limit. For example, with $n=2000$ and $\mu=0.04$, all five instances are solved to optimality by the direct approach. For those instances, the Benders approach is either able to prove optimality (for example, for $n=200$ and $\mu=0.04$, $O_{\text{Benders}}=0\%$) or terminate with a low optimality gap (for example, for $n=3000$ and $\mu=0.04$, $O_{\text{Benders}}=0.08\%$); among all instances with $\mu\in\{0.02,0.04\}$, the average optimality gap is no more than about 1.05%. More importantly, the solution obtained by the Benders approach is at least as good as the solution obtained after two hours of direct solution of SplitMIO, which can be seen from the fact that $G_{\text{Benders}}$ is 0.0% in all of these $(n,\mu)$ combinations.
The second type of behavior corresponds to “hard” instances, which are the instances with $\mu\in\{0.06,0.08,0.10,0.12\}$. For these instances, Gurobi generally struggles to solve the LO relaxation of SplitMIO within the two hour time limit. When this happens, the integer solution returned by Gurobi is obtained from applying heuristics before solving the root node of the branch-and-bound tree, which is often quite suboptimal. (This often results in integer solutions with an objective value of 0.0, leading to values of $O_{\text{Direct}}$ and $G_{\text{Direct}}$ that are close to 100%.) In contrast, the Benders approach delivers significantly better solutions; among all of these instances, $G_{\text{Benders}}$ is within 1%, indicating that on average the Benders approach is within 1% of the best solution of each instance. In addition, $O_{\text{Benders}}$ is generally lower than $O_{\text{Direct}}$, indicating that the Benders approach in general makes more progress towards proving optimality than the direct approach.
Overall, these results suggest that our Benders approach can deliver high quality solutions to large-scale instances of the assortment optimization problem in a reasonable computational timeframe that are at least as good, and often significantly better, than those obtained by the D&C heuristic or those obtained by directly solving the problem using Gurobi.
6 Numerical Experiments with Real Transaction Data
We examine the performance of the optimal assortment generated by the decision forest model on problem instances calibrated with a real-world transaction dataset. Different from Section 5, the problem instances in this section are of a smaller scale, which help us glean operational insights.
6.1 Background
We consider the IRI Academic Dataset (Bronnenberg et al. 2008), which is comprised of real-world transaction records of store sales for thirty product categories from forty-seven U.S. markets. The same dataset has been used by Jagabathula and Rusmevichientong (2019) to empirically demonstrate the loss of rationality in consumer choice and by Chen and Mišić (2022) to evaluate the predictive performance of the decision forest model.
We follow the literature to pre-process the raw transaction data. We first aggregate items with the same vendor code as a product, a common pre-processing technique in the marketing science community (Bronnenberg and Mela 2004, Nijs et al. 2007). Following the setup in the literature (Jagabathula and Rusmevichientong 2019, Chen and Mišić 2022) and focusing on the data from the first two weeks of the calendar year 2007 due to the data volume, we select the top nine purchased items as the products and combine the remaining items as the outside/no-purchase option. We further transform the transactions into assortment-choice pairs as follows. We first call $\mathcal{T}$ the set of transactions. For each transaction $\tau\in\mathcal{T}$, we have the following information: the week of the purchase $\tau_{\text{week}}$, the store where the transaction happened $\tau_{\text{store}}$, the sold product $\tau_{\text{prod}}$, and the selling price $\tau_{\text{price}}$. Let $\mathcal{W}$ and $\mathcal{Z}$ be the nonrepeated collection of $\{\tau_{\text{week}}\}_{\tau\in\mathcal{T}}$ and $\{\tau_{\text{store}}\}_{\tau\in\mathcal{T}}$, respectively. For each week $w\in\mathcal{W}$ and store $s\in\mathcal{Z}$, we define $S_{w,s}=\bigcup_{\tau\in\mathcal{T}}\{\tau_{\text{prod}}\mid\tau_{\text{week}}=w,\tau_{\text{store}}=s\}$ as the set of products that was purchased at least once at store $s$ during week $w$. The set of transactions $\mathcal{T}$ is thus transformed into the set of assortment-choice pairs as $\{\left(S_{\tau_{\text{week}},\tau_{\text{store}}},\tau_{\text{prod}}\right)\}_{\tau\in\mathcal{T}}$, which will be used for choice model estimation. We further define the marginal revenue $\rho_{i}$ as the average of the historical prices $\left(\tau_{\text{price}}\right)_{\tau\in\mathcal{T}:\tau_{\text{prod}}=i}$ for $i\in\mathcal{N}$.
6.2 Results: Improvement in Expected Revenue
We estimate both the ranking-based model and the decision forest model from the assortment-choice pairs $\{\left(S_{t_{\text{week}},t_{\text{store}}},t_{\text{prod}}\right)\}_{t\in\mathcal{T}}$ in each product category by maximum likelihood estimation. For details on estimating the ranking-based model, see van Ryzin and Vulcano (2014, 2017), and for the decision forest model, refer to Chen and Mišić (2022). In particular, we follow Chen and Mišić (2022) to warm start the solver by setting the initial solution as the estimated ranking-based model. For simplicity, we set the tree depth limit as three (i.e., leaves can have depth at most four). Typically, the tree depth is determined by cross validation. Chen and Mišić (2022) have reported that trees of depth three lead to good predictive performance, while deeper trees tend to overfit.
We further denote $S^{\text{RM}}$ and $S^{\text{DF}}$ as the optimal assortments under the estimated ranking-based model and the decision forest model, respectively. Note that we obtain $S^{\text{RM}}$ using an integer programming approach (Bertsimas and Mišić 2019, Feldman et al. 2019). We use $Z^{*}_{{\mathcal{C}}}$ to denote the maximal expected revenue under a given choice model $\mathcal{C}$ and $Z_{\mathcal{C}}(S)$ to denote the expected revenue of assortment $S$. Obviously, $Z^{*}_{\text{RM}}=Z_{\text{RM}}(S^{\text{RM}})$ and $Z^{*}_{\text{DF}}=Z_{\text{DF}}(S^{\text{DF}})$. To discuss the relative performance of assortments $S^{\text{RM}}$ and $S^{\text{DF}}$, we define the following two metrics: $RI_{\text{DF}}=100\%\times\frac{Z_{\text{DF}}(S^{\text{DF}})-Z_{\text{DF}}(S^{\text{RM}})}{Z_{\text{DF}}(S^{\text{RM}})}$ and $RI_{\text{RM}}=100\%\times\frac{Z_{\text{RM}}(S^{\text{RM}})-Z_{\text{RM}}(S^{\text{DF}})}{Z_{\text{RM}}(S^{\text{DF}})}$. The two metrics measure the relative improvement of the optimal assortment $S^{\text{DF}}$ (or $S^{\text{RM}}$) from $S^{\text{RM}}$ (or $S^{\text{DF}}$) under the decision forest model (or the ranking-based model). Note that both metrics are non-negative by their definitions. We also define the Hamming distance $\Delta_{\text{H}}=\sum_{i=1}^{n}|\mathbb{I}\left[i\in S^{\text{DF}}\right]-\mathbb{I}\left[i\in S^{\text{RM}}\right]|$ to measure how different the two assortments are.
Table 3 summarizes the comparison of $S^{\text{DF}}$ and $S^{\text{RM}}$ in terms of expected revenue under the measures $RI_{\text{DF}}$ and $RI_{\text{RM}}$. When the ground truth is the estimated decision forest model, the assortment $S^{\text{DF}}$ can outperform $S^{\text{RM}}$ up to $32\%$ and with $7\%$ improvement on average in the expected revenue. Meanwhile, when the ground truth is the estimated ranking-based model, the assortment $S^{\text{DF}}$ only performs $3\%$ worse than the optimal assortment $S^{\text{RM}}$ on average. Recall that the class of the ranking-based models is equivalent to the class of choice models of random utility maximization (RUM) principle, or, the class of rational choice models (Jagabathula and Rusmevichientong 2019). Table 3 suggests that the optimal assortment generated by the decision forest model can be quite beneficial when customer choice deviates from the rational choice theory, and does not lose much even if customers are strictly rational. We also remark that on average, the two assortments $S^{\text{DF}}$ and $S^{\text{RM}}$ only differ from each other with Hamming distance $1.9$, while they have similar sizes: the average sizes of $S^{\text{DF}}$ and $S^{\text{RM}}$ are 4.9 and 5.3, respectively.
Note that, in this real-data setting, the problem instances involve significantly fewer products compared to the synthetic-data setting described in Section 5. As a result, the optimal assortments $S^{\text{DF}}$ and $S^{\text{RM}}$ are obtained almost instantaneously (in less than one second) using the integer programming approach.
Category $RI_{\text{DF}}$ $RI_{\text{RM}}$ $\Delta_{\text{H}}$ Category $RI_{\text{DF}}$ $RI_{\text{RM}}$ $\Delta_{\text{H}}$
Beer 32.8 5.9 3 Mayonnaise 0.0 0.0 0
Blades 0.0 0.0 0 Milk 5.2 0.7 1
Carbonated Beverages 4.8 3.8 4 Mustard / Ketchup 5.6 1.1 1
Cigarettes 2.5 8.3 1 Paper Towels 1.0 2.4 2
Coffee 26.3 11.0 4 Peanut Butter 10.9 5.6 3
Cold Cereal 6.1 0.3 2 Photo 0.0 0.0 0
Deodorant 3.7 2.2 3 Salty Snacks 12.0 3.7 2
Diapers 0.0 0.0 0 Shampoo 19.0 1.4 4
Facial Tissue 0.2 0.6 1 Soup 13.0 5.0 3
Frozen Dinners 2.7 5.2 2 Spaghetti / Italian Sauce 4.9 3.3 2
Frozen Pizza 10.9 4.0 4 Sugar Substitutes 10.7 6.0 1
Household Cleaners 16.0 6.6 4 Toilet Tissue 0.0 0.0 0
Hotdogs 18.7 1.4 4 Toothbrush 0.0 0.0 0
Laundry Detergent 0.1 1.9 1 Toothpaste 2.7 0.9 2
Margarine / Butter 5.3 12.9 1 Yogurt 3.8 2.1 1
Average over all categories 7.3 3.2 1.9
Table 3: The relative performance of the assortments generated by the decision forest model and the ranking-based model in each product category in the IRI Academic Dataset
6.3 Case Study: Beer Category
To investigate why we observed high relative improvement with the assortment $S^{\text{DF}}$ over $S^{\text{RM}}$ in some categories, we look into the Beer category, where the relative improvement is 32.8%.
We first observe that the consumer choice within the Beer category exhibits a notable phenomenon known as choice overload, a well-documented behavioral anomaly in the field of marketing science (Iyengar and Lepper 2000, Chernev et al. 2015, Long et al. 2023). This phenomenon describes the situation in which an abundance of available products can actually deter customers from making a purchase. Choice overload contradicts with the principles of rational choice theory. According to this theory, when a store increases the assortment size, a rational consumer should theoretically be more inclined to make a purchase from the assortment. This is because a larger assortment would increase the chance that the customer finds one product that aligns with her preference. In the context of choice modeling, a rational choice model satisfies the following:
$P(0\mid S)\geq P(0\mid S^{\prime})\quad\text{ whenever }\quad S\subseteq S^{\prime}.$ (23)
Note that equation (23) directly stems from the regularity property, which the ranking-based model and other RUM choice models, such as the mixed-MNL model, always adhere to (Rieskamp et al. 2006, Jagabathula and Rusmevichientong 2019). However, real-world scenarios often deviate from equation (23). When provided with more options, customers may become fatigued from the search process or experience a decrease in their confidence in decision-making, leading them to opt not to make a purchase. In the consumer psychology literature, Iyengar and Lepper (2000) conducted a field experiment demonstrating that individuals are more likely to purchase gourmet jams or chocolates when presented with a smaller assortment of size six rather than a more extensive assortment of size 24 or 30.
Figure 3: (left) The choice overload effect in the Beer category. Each scatter point represents the empirical no-purchase probability $\bar{P}(0\mid S)$ for an historical assortment $S\in\mathcal{S}_{\text{hist}}$; (right) The average no-purchase probability of the estimated ranking-based and decision forest models (with the error bars as the standard deviation). The black triangle represents that $P_{\text{DF}}(0\mid S^{\text{DF}})=8.6\%$
Figure 3 (left) visually depicts the prevalence of choice overload within the Beer category of the IRI dataset. Each data point in the figure corresponds to a pair ($|S|$, $\bar{P}(0\mid S)$), where $|S|$ represents the assortment size of a historical assortment $S\in\mathcal{S}_{\text{hist}}\equiv\{S_{w,s}\mid w\in\mathcal{W},s\in\mathcal{S}\}$, and $\bar{P}(0\mid S)$ represents the empirical probability of the no-purchase option being chosen when the assortment $S$ was offered. For a better illustration, we have excluded assortments with very few transactions (less than or equal to 10) from Figure 3 (left). In this figure, the rightmost data point, denoted by a red triangle, corresponds to the no-purchase probability when the assortment contains all available products ($|S|=9$). It is important to note that in a scenario where equation (23) holds, indicating the absence of choice overload, each data point should exhibit a higher $y$-value than the red triangle. However, this condition is met for only 8 out of the 41 data points in the figure, suggesting a violation of equation (23) within the Beer category.
The presence of choice overload in the data has a significant impact on the estimation outcomes of both the ranking-based model and the decision forest model, leading to divergent assortment decisions. To begin, we define $\bar{P}_{\mathcal{C}}(0\mid m)$ as the average no-purchase probability when presented with an assortment of size $m$; formally, $\bar{P}_{\mathcal{C}}(0\mid m)={\sum_{S\subseteq\mathcal{N}\,:\,|S|=m}P_{\mathcal{C}}(0\mid S)}\big{/}{|\{S\subseteq\mathcal{N}\,:\,|S|=m\}|}$, where $P_{\mathcal{C}}(0\mid S)$ represents the no-purchase probability given assortment $S$ under a choice model $\mathcal{C}$. We consider two choice models, the decision forest model (DF) and the ranking-based model (RM), both estimated from data. Figure 3 (right) presents the average no-purchase probability curves, $\bar{P}_{\mathcal{C}}(0\mid m)$, for these two models, with the error bars representing the standard deviation. Recall that the ranking-based model adheres to the regularity property and strictly follows equation (23). Consequently, the resulting no-purchase probability curve (depicted in blue) does not exhibit the choice overload effect; it consistently decreases as we expand the assortment size. In contrast, the decision forest model is not bound by the regularity property and has the capacity to learn the choice overload effect from the data. The resulting no-purchase probability curve (shown in red) indicates that customers are indeed more inclined to make a purchase as the assortment initially expands. However, after reaching an assortment size of $|S|=6$, the choice overload effect becomes evident. Customers become less motivated, and we observe an increase in the no-purchase probability when $|S|\geq 7$.
The extent to which the decision forest model and the ranking-based model capture the phenomenon of choice overload from the data significantly influences their downstream assortment decisions. It is important to note that within the dataset, the marginal revenues $\rho_{i}$ of the products exhibit low variation, ranging between $8.14$ and $10.83$ USD (keeping in mind that beer is often sold in packs of six). Since prices are quite uniform and well above zero, the optimal strategy is not about making customers choose a specific product but rather about making them buy any product from the offered assortment, i.e., reducing the no-purchase probability.
For the ranking-based model, the optimal assortment $S^{\text{RM}}$ includes all products, resulting in $|S^{\text{RM}}|=9$, since this is the only way it can minimize the no-purchase probability. In contrast, the decision forest model, having captured the choice overload effect from the data, takes a different approach. It examines assortments of size six and identifies one that decreases the no-purchase probability from ${P}_{\text{DF}}(0\mid S^{\text{RM}})=33.1\%$ to ${P}_{\text{DF}}(0\mid S^{\text{DF}})=8.6\%$ (see the black triangle in Figure 3 (right)). This strategic choice results in a remarkable 32% improvement in expected revenue, measured according to the decision forest model. This highlights the advantage of the assortment planning approach proposed in this paper: by leveraging the decision forest model’s ability to capture consumer choice patterns, businesses can tailor their product offerings to effectively capitalize on consumers’ departures from strictly rational purchasing behavior.
7 Conclusions
We developed a mixed-integer optimization methodology for solving the assortment optimization problem under the decision forest model, which accounts for both rational and non-rational customer behaviors. We first established the inapproximability of the problem, then presented two formulations, linking them to the optimization models in the literature. To enhance scalability, we proposed a large-scale solution approach based on Benders decomposition. Through synthetic data experiments, we demonstrated the efficiency of our methods, and with real-world data, we showed how consumer behavioral anomalies can be learned and leveraged within our framework.
Acknowledgments
We sincerely thank the department editor Huseyin Topaloglu, the associate editor and two anonymous reviewers for their thoughtful comments that have helped to significantly improve the paper.
[…]